# Surface decomposition method for sensitivity analysis of first-passage dynamic reliability of linear systems

### v1

@@ 4 Numerical examples
In this section, two numerical examples, including an oscillator and a shear-type structure, are presented to demonstrate the effectiveness of the proposed surface decomposition method (SDM) for sensitivity analysis of first-passage dynamic reliability of linear systems. The first-passage dynamic reliability analysis of the linear system is also conducted by the efficient importance sampling (ISEE) au2001first . The finite difference method in conjunction with importance sampling, which can be denoted as FDM-IS, is adopted to obtain the reference solutions to the sensitivities of failure probabilities. The basic idea of ISEE and FDM-IS are presented in A and B, respectively.
@@ 4.1 An oscillator
The circular frequency $\omega_{n}$ and damping ratio $\zeta_{n}$ of the oscillator are selected as the design parameters. To investigate the effect of the failure probability magnitude on the efficiency of the proposed method for sensitivity analysis of dynamic reliability, four cases are considered for the symmetric threshold, i.e., $c=0.013\mathrm{m}$, $c=0.016\mathrm{m}$, $c=0.018\mathrm{m}$, and $c=0.020\mathrm{m}$. The results of failure probabilities for these cases, obtained from ISEE with a target coefficient of variation (COV) of 0.1, are presented in Table 1. The corresponding sensitivities of the failure probability with respect to $\omega_{n}$ and $\zeta_{n}$, computed by the proposed SDM with a target COV of 0.1, are shown in Table 2, together with the reference solutions obtained from FDM-IS using a 0.1% perturbation in the design parameters and a target COV of 0.02. The convergence histories of the sensitivities of failure probability, evaluated by SDM for all cases, are depicted in Figure 2.
@@
Case $c$ (m) $\mathbb{P}$ Function Evaluations COV
1 0.013 $3.06\times 10^{-3}$ 35 0.099
2 0.016 $1.92\times 10^{-5}$ 30 0.097
3 0.018 $3.85\times 10^{-7}$ 23 0.098
4 0.020 $4.01\times 10^{-9}$ 20 0.096
@@
1 0.013 SDM $7.37\times 10^{-3}$ 0.100 $5.95\times 10^{-1}$ 0.046 839
@@
2 0.016 SDM $6.86\times 10^{-5}$ 0.100 $5.62\times 10^{-3}$ 0.054 437
@@
3 0.018 SDM $1.73\times 10^{-6}$ 0.100 $1.44\times 10^{-4}$ 0.050 346
@@
4 0.020 SDM $2.46\times 10^{-8}$ 0.100 $1.98\times 10^{-6}$ 0.053 267
@@
Figure 2: Convergence histories of sensitivities of failure probability with respect to $\omega_{n}$ and $\zeta_{n}$ for Example 1. (a) Case 1: $c=0.013\mathrm{m}$; (b) Case 2: $c=0.016\mathrm{m}$; (c) Case 3: $c=0.018\mathrm{m}$; (d) Case 4: $c=0.020\mathrm{m}$. The number of function evaluations required in the proposed SDM for Case 1, Case 2, Case 3 and Case 4 are 839, 437, 346 and 267, respectively.
It can be observed from Table 1 that ISEE exhibits very high efficiency in computing the failure probabilities, with its efficiency increasing as the failure probability decreases. This phenomenon is consistent with the findings shown in au2001first , katafygiotis2006domain . Interestingly, the proposed SDM also possesses this counterintuitive characteristic in the context of sensitivity analysis, exhibiting higher efficiency as the failure probability becomes smaller, as shown in Table 2. Overall, the sensitivity results obtained from SDM are in very good agreement with the reference solutions obtained from FDM-IS, while requiring less than 1000 function evaluations, thereby demonstrating the effectiveness of the proposed method.
@@
Case 4 is taken for further illustration. The proposed SDM requires only 267 function evaluations to compute the sensitivities of failure probability with respect to $\omega_{n}$ and $\zeta_{n}$. In fact, if the sensitivity with respect to $\zeta_{n}$ alone is of interest, fewer function evaluations are required due to its faster convergence rate, as shown in Figure 2(d), where less than 100 function evaluations are sufficient to meet the stopping criterion. For the reference method FDM-IS, $1.36\times 10^{5}$ and $3.47\times 10^{5}$ function evaluations are needed to obtain the sensitivity results with respect to $\omega_{n}$ and $\zeta_{n}$, respectively.
@@ 4.2 A shear-type structure
The stiffness and damping coefficients of the viscoelastic dampers, $k_{ve,k}$ and $c_{ve,k}\;(k=1,2,\dots,20)$, are selected as the design parameters. Three cases are considered for the intensity parameter, i.e., $S_{0}=0.010\mathrm{m^{2}/s^{3}}$, $S_{0}=0.008\mathrm{m^{2}/s^{3}}$ and $S_{0}=0.007\mathrm{m^{2}/s^{3}}$, which leads to different magnitudes of the failure probability. The results of failure probabilities are presented in Table 3, obtained from ISEE using 188, 142 and 95 function evaluations for Case 1, Case 2 and Case 3, respectively. The results of sensitivities of failure probability with respect to all the design parameters $k_{ve,k}$ and $c_{ve,k}\;(k=1,2,\dots,20)$ are depicted in Figure 4, which are obtained from the proposed SDM using 1663, 953 and 877 function evaluations for Case 1, Case 2 and Case 3, respectively. The target COV is set to be 0.1 for both ISEE and SDM. Consistent with the observation in Example 1, the ISEE and the proposed SDM exhibit higher computational efficiency as the failure probability decreases. It is seen from Figure 4 that the failure probability of the structure is most sensitive to the parameters of the first-story viscoelastic damper, i.e., $k_{ve,1}$ and $c_{ve,1}$, while the sensitivity computation with respect to the damping coefficient of the top-story viscoelastic damper, i.e., $c_{ve,20}$, dominates the convergence behavior of SDM.
@@
Case $S_{0}\;(\mathrm{m^{2}/s^{3}})$ $\mathbb{P}$ Function Evaluations COV
1 0.010 $3.83\times 10^{-3}$ 188 0.100
2 0.008 $3.63\times 10^{-4}$ 142 0.099
3 0.007 $7.04\times 10^{-5}$ 95 0.099
@@
The sensitivities of the failure probability with respect to $k_{ve,1}$ and $c_{ve,1}$ are further investigated, as summarized in Table 4. The reference solutions obtained from FDM-IS, with a target COV of 0.1 and a 0.1% perturbation of design parameter, are also provided for comparison. The convergence histories of these sensitivities are illustrated in Figure 5. It is seen from Table 4 that the sensitivity results obtained from the proposed SDM align very well with the reference solutions. It is noted that the number of function evaluations for SDM reported in Table 4 corresponds to the sensitivity computation with respect to all 40 design parameters. In fact, if only the sensitivities with respect to $k_{ve,1}$ and $c_{ve,1}$ are of interest, even fewer function evaluations are required to attain the target COV of 0.1, which can be observed from Figure 5. Overall, the proposed SDM demonstrates excellent accuracy and efficiency for the sensitivity analysis of first-passage dynamic reliability of linear systems.
@@
1 0.010 SDM $1.06\times 10^{-9}$ 0.070 $3.93\times 10^{-9}$ 0.071 1663
@@
2 0.008 SDM $1.56\times 10^{-10}$ 0.075 $5.75\times 10^{-10}$ 0.076 953
@@
3 0.007 SDM $3.10\times 10^{-11}$ 0.077 $1.16\times 10^{-10}$ 0.078 877
@@
Figure 5: Convergence histories of sensitivities of failure probability with respect to $k_{ve,1}$ and $c_{ve,1}$ for Example 2. (a) Case 1: $S_{0}=0.010\mathrm{m^{2}/s^{3}}$; (b) Case 2: $S_{0}=0.008\mathrm{m^{2}/s^{3}}$; (c) Case 3: $S_{0}=0.007\mathrm{m^{2}/s^{3}}$. The number of function evaluations required in the proposed SDM for Case 1, Case 2 and Case 3 are 1663, 953 and 877, respectively.

### v2

@@ 4 Numerical examples
In this section, three numerical examples, including an oscillator, a shear-type structure and a building frame structure, are investigated to demonstrate the effectiveness of the proposed surface decomposition method (SDM) for sensitivity analysis of first-passage dynamic reliability of linear systems. The first-passage dynamic reliability analysis of the linear system is also conducted, utilizing the efficient importance sampling (ISEE) au2001first, the domain decomposition method (DDM) katafygiotis2006domain and the directional importance sampling (DIS) misraji2020application. The DIS is also employed to perform the reliability sensitivity analysis valdebenito2021sensitivity. The finite difference method in conjunction with the importance sampling, which can be denoted as FDM-IS, is adopted to obtain the reference solutions to the sensitivities of failure probabilities. The basic idea of ISEE and FDM-IS are presented in A and B, respectively.
@@ 4.1 An oscillator
The circular frequency $\omega_{n}$ and damping ratio $\zeta_{n}$ of the oscillator are selected as the design parameters. To investigate the effect of the failure probability magnitude on the efficiency of the proposed method for sensitivity analysis of dynamic reliability, four cases are considered for the symmetric threshold, i.e., $c=0.013\mathrm{m}$, $c=0.016\mathrm{m}$, $c=0.018\mathrm{m}$, and $c=0.020\mathrm{m}$. The results of failure probabilities for these four cases, obtained from ISEE, DDM and DIS with a target coefficient of variation (COV) of 0.1, are presented in Table 1. The corresponding sensitivities of the failure probability with respect to $\omega_{n}$ and $\zeta_{n}$, computed by the proposed SDM and the DIS with a target COV of 0.1, are shown in Table 2, together with the reference solutions obtained from FDM-IS using a 0.1% perturbation in the design parameters and a target COV of 0.02. It is noted that, since both SDM and DIS are able to reuse the results of function evaluations when estimating the sensitivities of failure probability with respect to different design parameters, only a single number of function evaluations is reported for SDM and DIS in Table 2. However, for FDM-IS, two separate numbers of function evaluations are given for the sensitivity estimation with respect to $\omega_{n}$ and $\zeta_{n}$, respectively. The convergence histories of the sensitivities of failure probability, evaluated by SDM and DIS for all cases, together with the corresponding COVs and the relative errors with respect to the reference solutions, are depicted in Figure 2.
@@
Case $c$ (m) Method $\mathbb{P}$ Function Evaluations COV
1 0.013 ISEE $3.06\times 10^{-3}$ 35 0.099
DDM $3.32\times 10^{-3}$ 32 0.099
DIS $3.28\times 10^{-3}$ 28 0.097
2 0.016 ISEE $1.92\times 10^{-5}$ 30 0.097
DDM $2.06\times 10^{-5}$ 29 0.099
DIS $1.94\times 10^{-5}$ 21 0.099
3 0.018 ISEE $3.85\times 10^{-7}$ 23 0.098
DDM $3.67\times 10^{-7}$ 24 0.099
DIS $3.63\times 10^{-7}$ 15 0.093
4 0.020 ISEE $4.01\times 10^{-9}$ 20 0.096
DDM $4.04\times 10^{-9}$ 18 0.099
DIS $4.26\times 10^{-9}$ 11 0.081
@@
1 0.013 SDM $7.30\times 10^{-3}$ 0.100 $6.14\times 10^{-1}$ 0.049 714
DIS $6.32\times 10^{-3}$ 0.100 $5.24\times 10^{-1}$ 0.054 1673
@@
2 0.016 SDM $6.80\times 10^{-5}$ 0.100 $5.67\times 10^{-3}$ 0.049 501
DIS $6.96\times 10^{-5}$ 0.100 $5.70\times 10^{-3}$ 0.051 751
@@
3 0.018 SDM $1.62\times 10^{-6}$ 0.100 $1.42\times 10^{-4}$ 0.052 326
DIS $1.58\times 10^{-6}$ 0.100 $1.43\times 10^{-4}$ 0.048 615
@@
4 0.020 SDM $2.40\times 10^{-8}$ 0.100 $2.02\times 10^{-6}$ 0.053 252
DIS $2.25\times 10^{-8}$ 0.100 $1.92\times 10^{-6}$ 0.058 365
@@
Figure 2: Convergence histories of sensitivities of failure probability with respect to $\omega_{n}$ and $\zeta_{n}$ for Example 1. (a) Case 1: $c=0.013\mathrm{m}$; (b) Case 2: $c=0.016\mathrm{m}$; (c) Case 3: $c=0.018\mathrm{m}$; (d) Case 4: $c=0.020\mathrm{m}$. The number of function evaluations required in the proposed SDM for Case 1, Case 2, Case 3 and Case 4 are 714, 501, 326 and 252, respectively, while those required in DIS are 1673, 751, 615, 365, respectively. Accordingly, the speedup ratios of the proposed SDM over DIS for these four cases are approximately 2.34, 1.50, 1.89 and 1.45, respectively.
It can be observed from Table 1 that all three methods, namely ISEE, DDM and DIS, exhibit very high efficiency in computing the failure probabilities, with their efficiency increasing as the failure probability decreases. This phenomenon is consistent with the findings presented in au2001first, katafygiotis2006domain. Interestingly, the proposed SDM and the DIS also possess this counterintuitive characteristic in the context of reliability sensitivity analysis, demonstrating higher efficiency as the failure probability becomes smaller, as shown in Table 2. The mechanism underlying the above phenomena is that the efficiency of all these methods, whether applied to first-passage dynamic reliability analysis or reliability sensitivity analysis, is strongly dependent on the degree of overlap among the component failure domains/surfaces. A lower overlapping degree among the component failure domains/surfaces corresponds to higher efficiency of these methods. In general, a first-passage failure event with a lower probability implies larger distances between the origin and the component failure domains/surfaces, which results in a lower degree of overlap and finally leads to higher efficiency of these methods.
@@
Overall, the sensitivity results obtained from SDM and DIS are in good agreement, and they also agree well with the reference solutions obtained from FDM-IS. Moreover, the SDM exhibits higher efficiency than DIS, requiring less than 1000 function evaluations for all cases, thereby demonstrating the effectiveness of the proposed method. Case 1 is taken for further illustration. The proposed SDM requires only 714 function evaluations to compute the sensitivities of failure probability with respect to $\omega_{n}$ and $\zeta_{n}$, while 1673 function evaluations are required in DIS meet the same stopping criterion. The speedup ratio of SDM over DIS is approximately 2.34. For the reference method FDM-IS, $1.36\times 10^{5}$ and $3.47\times 10^{5}$ function evaluations are needed to obtain the sensitivity results with respect to $\omega_{n}$ and $\zeta_{n}$, respectively.
In the present SDM, the importance sampling PMF is chosen as $h_{1}(i)=\mathbb{P}_{i}/\sum_{s=1}^{n}\mathbb{P}_{s}$, where $\mathbb{P}_{i}\;(i=1,2,\dots,n)$ are the component failure probabilities. To demonstrate that this PMF provides a sufficient surrogate for the theoretically optimal PMF, the optimal PMFs with respect to $\omega_{n}$ and $\zeta_{n}$, denoted by $h_{2}(i)=\eta_{i}[\theta=\omega_{n}]/\sum_{s=1}^{n}\eta_{s}[\theta=\omega_{n}]$ and $h_{3}(i)=\eta_{i}[\theta=\zeta_{n}]/\sum_{s=1}^{n}\eta_{s}[\theta=\zeta_{n}]$, respectively, are also evaluated. Here, $\eta_{i}[\theta=\omega_{n}]$ and $\eta_{i}[\theta=\zeta_{n}]\;(i=1,2,\dots,n)$ denote the component surface integrals with respect to $\omega_{n}$ and $\zeta_{n}$, respectively, which are estimated by the direct Monte Carlo simulation, and the corresponding results are presented in Figures 3(a) and 3(b). The resulting optimal PMFs $h_{2}(i)$ and $h_{3}(i)$ as well as the adopted PMF $h_{1}(i)$ are depicted in Figure 3(c), from which it can be seen that $h_{1}(i)$ serves as a good surrogate for the optimal PMFs. Notably, the fact that sensitivity estimations with respect to different design parameters share the same surrogate PMF is the key reason that enables the reuse of function evaluation results in SDM.
Figure 3: Comparison among different PMFs. (a) Estimation of component surface integrals with respect to $\omega_{n}$; (b) Estimation of component surface integrals with respect to $\zeta_{n}$; (c) Comparison among different PMFs.
To investigate the computational efficiency of SDM for higher input dimensionality and larger number of component failure events, the SDM is further employed to perform the reliability sensitivity analysis for different input dimensionalities $d$ and different numbers of component failure events $n$. Two scenarios are considered: (a) $d$ varies from 1000 to 10000 with $n=1000$ fixed; (b) $n$ varies from 1000 to 10000 with $d=1000$ fixed. Different values of $d$ and $n$ can be achieved by scaling the circular frequency interval $\Delta\omega$ and time duration $T$, respectively. The number of function evaluations required in SDM corresponding to the same target COV of 0.1 for these two scenarios are shown in Figures 4(a) and 4(b), respectively. It can be seen from Figure 4 that the efficiency of SDM is independent of the input dimensionality, while it decreases as the number of component failure events increases. This is expected since a larger number of samples are required to adequately account for the contributions from the additional component surface integrals to the total sensitivity estimation. It should be noted that, in practical engineering scenarios with nonstationary excitations, an increase in the number of component failure events does not necessarily imply a reduction in SDM efficiency, since only the component surface integrals corresponding to the strong-motion intervals contribute significantly to the total sensitivity estimation.
Figure 4: Comparison of computational efficiency of SDM for different input dimensionalities $d$ and numbers of component failure events $n$. (a) Number of function evaluations versus $d$, with $d\in[1000,10000]$ and fixed $n=1000$; (b) Number of function evaluations versus $n$, with $n\in[1000,10000]$ and fixed $d=1000$.
@@ 4.2 A shear-type structure
The stiffness and damping coefficients of the viscoelastic dampers, $k_{ve,k}$ and $c_{ve,k}\;(k=1,2,\dots,20)$, are selected as the design parameters. Three cases are considered for the intensity parameter, i.e., $S_{0}=0.010\mathrm{m^{2}/s^{3}}$, $S_{0}=0.008\mathrm{m^{2}/s^{3}}$ and $S_{0}=0.007\mathrm{m^{2}/s^{3}}$, which leads to different magnitudes of the failure probability. The results of failure probabilities are presented in Table 3, from which it can be observed that all three methods, including ISEE, DDM and DIS, exhibit very high efficiency for the first-passage dynamic reliability analysis, and DIS is the most efficient among them. The results of sensitivities of failure probability with respect to all design parameters $k_{ve,k}$ and $c_{ve,k}\;(k=1,2,\dots,20)$ are depicted in Figure 6, which are obtained from the proposed SDM with a target COV of 0.1 using 1663, 953 and 877 function evaluations for Case 1, Case 2 and Case 3, respectively. Consistent with the observation in Example 1, SDM exhibits higher computational efficiency as the failure probability decreases. It is seen from Figure 6 that the failure probability of the structure is most sensitive to the parameters of the first-story viscoelastic damper, i.e., $k_{ve,1}$ and $c_{ve,1}$, while the sensitivity computation with respect to the damping coefficient of the top-story viscoelastic damper, i.e., $c_{ve,20}$, dominates the convergence behavior of SDM.
@@
Case $S_{0}$ (m2/s3) Method $\mathbb{P}$ Function Evaluations COV
1 0.010 ISEE $3.83\times 10^{-3}$ 188 0.100
DDM $3.79\times 10^{-3}$ 183 0.100
DIS $3.49\times 10^{-3}$ 100 0.100
2 0.008 ISEE $3.63\times 10^{-4}$ 142 0.099
DDM $3.56\times 10^{-4}$ 139 0.099
DIS $3.82\times 10^{-4}$ 79 0.098
3 0.007 ISEE $7.04\times 10^{-5}$ 95 0.099
DDM $6.77\times 10^{-5}$ 84 0.099
DIS $72\times 10^{-5} 66$ 0.099
@@
The sensitivities of the failure probability with respect to $k_{ve,1}$ and $c_{ve,1}$ are further investigated using SDM and DIS with a target COV of 0.1, as summarized in Table 4. The reference solutions obtained from FDM-IS, with a target COV of 0.1 and a 0.1% perturbation of design parameter, are also provided for comparison. The convergence histories of these sensitivities as well as the corresponding COVs and the errors relative to the reference solutions are illustrated in Figure 7. It can be observed that the sensitivity results obtained from the proposed SDM and the DIS are in good agreement and both align well with the reference solutions. As for the efficiency, SDM outperforms DIS with speedup ratios of about 1.86, 1.65 and 1.29 for Case 1, Case 2 and Case 3, respectively. Overall, the proposed SDM demonstrates excellent accuracy and efficiency for the sensitivity analysis of first-passage dynamic reliability of linear systems.
@@
1 0.010 SDM $1.10\times 10^{-9}$ 0.099 $4.17\times 10^{-9}$ 0.100 782
DIS $0.92\times 10^{-9}$ 0.100 $3.53\times 10^{-9}$ 0.097 1458
@@
2 0.008 SDM $1.54\times 10^{-10}$ 0.098 $5.70\times 10^{-10}$ 0.100 558
DIS $1.46\times 10^{-10}$ 0.100 $5.46\times 10^{-10}$ 0.097 920
@@
3 0.007 SDM $3.08\times 10^{-11}$ 0.098 $1.15\times 10^{-10}$ 0.100 540
DIS $2.77\times 10^{-11}$ 0.100 $1.06\times 10^{-10}$ 0.100 697
@@
Figure 7: Convergence histories of sensitivities of failure probability with respect to $k_{ve,1}$ and $c_{ve,1}$ for Example 2. (a) Case 1: $S_{0}=0.010\mathrm{m^{2}/s^{3}}$; (b) Case 2: $S_{0}=0.008\mathrm{m^{2}/s^{3}}$; (c) Case 3: $S_{0}=0.007\mathrm{m^{2}/s^{3}}$. The number of function evaluations required in the proposed SDM for Case 1, Case 2 and Case 3 are 782, 558 and 540, respectively, whereas those required in DIS are 1458, 920 and 697, respectively. The speedup ratios of SDM over DIS for these three cases are approximately 1.86, 1.65 and 1.29, respectively.
4.3 A building frame structure
Consider a 4-storey reinforced concrete building frame structure with a storey height of 4.5m, as shown in Figure 8. The two planar frames and located on both sides of the building each consist of 7 bays with a bay width of 4.9m. The columns and beams for the frame structure are designed with rectangular cross sections measuring $0.4\mathrm{m}\times 0.5\mathrm{m}$ and $0.8\mathrm{m}\times 0.25\mathrm{m}$, respectively. The damping matrix is defined using the Rayleigh damping model, with a critical damping ratio of 0.05 assigned to the first and fifteenth modes of the structure. To suppress the excessive displacements of the structure under the Y-direction seismic excitation, each of the planar frames and is equipped with 28 linear viscous dampers, as illustrated in Figure 9. The dampers in frames and are characterized by the damping coefficients $c_{v,1},c_{v,2},\dots,c_{v,28}$ and $c_{v,29},c_{v,30},\dots,c_{v,56}$, respectively. The angle between the braced arm of the viscous damper and the floor is $\alpha=\arctan(4.5/4.9)$. The finite element model of the structure is developed in SAP2000, in which 1201 frame elements are used to model the columns and beams, and 530 shell elements are employed to simulate the floor slabs, resulting in a total of 4020 degrees of freedom.
Figure 8: A 4-storey reinforced concrete building frame structure.
Figure 9: Layouts of linear viscous dampers for planar frames and .
The structure is subjected to a uniformly modulated non-stationary Gaussian ground motion acceleration $\ddot{u}_{g}(\boldsymbol{X},t)=g(t)\ddot{u}_{g}^{0}(\boldsymbol{X},t)$, in which the intensity modulation function $g(t)$ and the correlation function of the stationary part $\ddot{u}_{g}^{0}(\boldsymbol{X},t)$ are the same as those defined in Eqs.(43) and (44), respectively, except that the intensity parameter is specified as $S_{0}=3.5\times 10^{-3}\mathrm{m}^{2}/\mathrm{s}^{3}$ in this example. The orthogonal decomposition method is also adopted to represent the seismic excitation, in which the time step and time duration are taken as $\Delta t=0.02\mathrm{s}$ and $T=30\mathrm{s}$, respectively, leading to $n=1500$ time steps. Therefore, the dimension of the standard Gaussian random vector $\boldsymbol{X}$ is 1500.
The failure event $F$ is defined as the exceedance of a prescribed symmetric threshold $c=0.01\mathrm{m}$ by the maximum absolute displacement among the 4 corner points A, B, C and D (see Figure 9) over a time duration of $T=30\mathrm{s}$. The failure event $F$ can be expressed as
$F=\left\{\,c-\max_{i\in\{\mathrm{A,B,C,D}\}}\max_{t\in[0,T]}|u_{i}(\boldsymbol{X},t)|\leq 0\,\right\}$ (48)
where $u_{i}(\boldsymbol{X},t)\;(i\in\{\mathrm{A,B,C,D}\})$ are the Y-direction horizontal displacements at points A, B, C and D, respectively. In this example, since the number of time steps is $n=1500$ and the number of critical response components is $m=4$, a total of $mn=6000$ component failure events are considered.
The damping coefficients of all the linear viscous dampers are selected as the design parameters, and they are set as $c_{v,k}=200\mathrm{kN\cdot s/m}\;(k=1,2,\dots,56)$. At the initial setup stage of SDM, the explicit expressions of the critical responses as well as their sensitivities with respect to all design parameters $c_{v,k}\;(k=1,2,\dots,56)$ need to be constructed. In principle, this would require a response time-history analysis plus 56 response sensitivity time-history analyses of the building frame structure. To reduce the computational cost, the adjoint variable method developed in hu2016explicit is employed, where the 56 response sensitivity time-history analyses are replaced by solving the adjoint equations corresponding to the 4 critical responses. This is equivalent to performing 4 response time-history analyses of the structure. It should be noted that the initial setup cost of SDM, which now amounts to 5 response time-history analyses of the structure, remains almost unchanged even when additional design parameters are introduced.
The results of failure probabilities obtained from ISEE, DDM and DIS are presented in Table 5, from which it can be observed that all three methods demonstrate very high efficiency for computing the first-passage failure probability, with DIS achieving the best performance. The results of the failure probability sensitivities with respect to all design parameters $c_{v,k}\;(k=1,2,\dots,56)$, which are obtained through the proposed SDM using 502 function evaluations with a target COV of 0.1, are illustrated in Figure 10. It can be observed that the failure probability sensitivities with respect to $c_{v,k}\;(k=1,2,\dots,28)$ are positive, while those with respect to $c_{v,k}\;(k=29,30,\dots,56)$ are negative, meaning that increasing the damping coefficients of the dampers installed on frame leads to an increase in the failure probability, whereas increasing the damping coefficients of the dampers on frame reduces the failure probability. This is reasonable when the system failure is governed by the displacement responses at points C and D on frame . In this case, increasing the damping on frame and frame may result in completely different effects on the displacement responses of frame , thereby leading to opposite trends in the failure probability.
It can be also observed from Figure 10 that the failure probability is more sensitive to the damping coefficients of the dampers on the first storey than to those on the upper storeys. Furthermore, the failure probability exhibits the highest positive sensitivity to the damping coefficient $c_{v,1}$ and the highest negative sensitivity to the damping coefficient $c_{v,29}$. With the results of failure probability and the corresponding sensitivities at hand, certain gradient-based optimizers, such as the method of moving asymptotes svanberg1987method, can be employed to update the design parameters $c_{v,k}\;(k=1,2,\dots,56)$. This process is typically repeated within the framework of reliability-based design optimization until convergence of the design parameters is achieved.
Table 5: Results of failure probabilities for Example 3.
Method $\mathbb{P}$ Function Evaluations COV
ISEE $1.76\times 10^{-3}$ 94 0.099
DDM $1.73\times 10^{-3}$ 84 0.099
DIS $1.74\times 10^{-3}$ 59 0.099
Figure 10: Results of sensitivities of failure probability with respect to $c_{v,k}\;(k=1,2,\dots,56)$ for Example 3. The number of function evaluations required in the proposed SDM is 502.
The sensitivities of the failure probability with respect to $c_{v,1}$ and $c_{v,29}$ are further explored using SDM, DIS and FDM-IS (served as reference solutions) with a target COV of 0.1. The results are presented in Table 6, while the convergence histories of the failure probability sensitivities, along with the corresponding COVs and errors relative to the reference solutions, are depicted in Figure 11. It can be observed that both SDM and DIS exhibit high accuracy, while SDM outperforms DIS in terms of efficiency with a speedup ratio of 2.56. Overall, the proposed SDM demonstrates its potential for solving practical engineering problems.
Table 6: Results of sensitivities of failure probability with respect to $c_{v,1}$ and $c_{v,29}$ for Example 3.
Method $\partial\mathbb{P}/\partial c_{v,1}$ COV $\partial\mathbb{P}/\partial c_{v,29}$ COV Function Evaluations
SDM $9.68\times 10^{-10}$ 0.100 $8.68\times 10^{-9}$ 0.085 456
DIS $8.30\times 10^{-10}$ 0.100 $8.90\times 10^{-9}$ 0.083 1166
FDM-IS $9.54\times 10^{-10}$ 0.100 $8.55\times 10^{-9}$ 0.100 $8.93\times 10^{6}/2.12\times 10^{6}$
Figure 11: Convergence histories of sensitivities of failure probability with respect to $c_{v,1}$ and $c_{v,29}$ for Example 3. The number of function evaluations required in the proposed SDM and the DIS are 456 and 1166, respectively. The speedup ratio of SDM over DIS is approximately 2.56.